Annie is 3/4 of Sally's age. In 10 years, the sum of their ages will be 83. Find their present ages. Thank you!
Answer by amarjeeth123(570)

Let Sally's age be x.
Annie's age is 3/4x.
In 10 years their ages are (x+10) and (3/4x+10).
(x+10)+(3/4x+10)=83
x+3/4x=63
7/4x=63
x=4/7*63=36
Sally's age is 36 and Annie's age is 27.